OpenStudy (oswaldmurphy):

An expression is shown below: square root of 50 plus square root of 2 Which statement is true about the expression?

OpenStudy (oswaldmurphy):

It is irrational and equal to 2 multiplied by the square root of 13. It is irrational and equal to 6 multiplied by the square root of 2. It is rational and equal to 2. It is rational and equal to 6.

OpenStudy (oswaldmurphy):

@pooja195 @.Sam. @ganeshie8 @sammixboo

sam (.sam.):

I'll give you a hint \(\sqrt{a}+\sqrt{a}=2\sqrt{a}\)

sam (.sam.):

You can separate the \(\sqrt{50}\) as a start

OpenStudy (oswaldmurphy):

Would it be B?

OpenStudy (oswaldmurphy):

Thats what I got but didnt understand how...

sam (.sam.):

\[\sqrt{50}+\sqrt{2} \\ \\ =\sqrt{25 \times 2} + \sqrt{2} \\ \\ = \sqrt{5^2 \times 2} + \sqrt{2} \\ \\ =5\sqrt{2}+\sqrt{2} \\ \\ =?\]

OpenStudy (oswaldmurphy):

B. thanks1